OpenStudy (anonymous):

WHO'S THE SMARTEST!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!You want to start a lawn and land-scaping business. Your initial investment, including mowers, rakes, and other supplies, will cost $9500. For each lawn you mow, you project that you will spend $5.50 on gas as well as $7.50 in labor costs. You can charge $55 for each lawn. Let x be the number of lawns you mow. For this business venture: a. Write the cost function, C(x). b. Write the revenue function, R(x). c. Determine the break-even point. Describe what this means

OpenStudy (anonymous):

\[C(x)=9500+13x\] \[R(x)=55x\] set them equal to get the break even point

OpenStudy (anonymous):

55x=9500+13x 42x=9500 x=226 55(226)=12,430 breakeven, is 226,12,430